### v1

# UVLM: Benchmarking Video Language Model for Underwater World Understanding

Abstract
Recently, the remarkable success of large language models (LLMs) has achieved a profound impact on the field of artificial intelligence. Numerous advanced works based on LLMs have been proposed and applied in various scenarios. Among them, video language models (VidLMs) are particularly widely used. However, existing works primarily focus on terrestrial scenarios, overlooking the highly demanding application needs of underwater observation. To overcome this gap, we introduce UVLM, an under water observation benchmark which is build through a collaborative approach combining human expertise and AI models. To ensure data quality, we have conducted in-depth considerations from multiple perspectives. First, to address the unique challenges of underwater environments, we selected videos that represent typical underwater challenges including light variations, water turbidity, and diverse viewing angles to construct the dataset. Second, to ensure data diversity, the dataset covers a wide range of frame rates, resolutions, 419 classes of marine animals, and various static plants and terrains. Next, for task diversity, we adopted a structured design where observation targets are categorized into two major classes: biological and environmental. Each category includes content observation and change/action observation, totaling 20 distinct task types. Finally, we designed several challenging evaluation metrics to enable quantitative comparison and analysis of different methods. Experiments on two representative VidLMs demonstrate that fine-tuning VidLMs on UVLM significantly improves underwater world understanding while also showing potential for slight improvements on existing in-air VidLM benchmarks, such as VideoMME and Perception text. The dataset and prompt engineering will be released publicly.
1 Introduction
Figure 1: Challenges for VidLMs for understanding underwater videos. Our UVLM is proposed to overcome this gap and it enables the 7B VidLM to achieve performance comparable to closed-source models like GPT-4o and Gemini.
Video-language understanding [1, 2] stands at the forefront of multimedia research, empowering systems to interpret, reason about, and generate natural language descriptions of temporal and dynamic visual content [3]. Recent advances [3, 4, 5] in video-language models (VidLMs) have demonstrated impressive performance in tasks such as video captioning, temporal grounding, and visual question answering, focusing primarily on human-centric scenarios and common object interactions. Despite these achievements, a critical question remains: Can current VidLMs effectively understand videos captured in special imaging conditions, such as underwater environments?
This question is especially relevant as underwater environments constitute an uncharted domain with immense scientific value (e.g., marine biodiversity monitoring, ecosystem health assessment) and substantial engineering applications (e.g., autonomous underwater vehicles, offshore infrastructure inspection). As illustrated in Figure 1, applying VidLMs to underwater content presents three critical challenges that distinguish this domain from conventional video-language tasks:
  • Degraded Visual Features Hindering Analytical Decisions. The underwater domain introduces fundamental perceptual barriers that confound existing video understanding approaches. Underwater scenes exhibit variable illumination with rapid light attenuation at depth, wavelength-dependent color distortion that shifts the visual spectrum, fluctuating turbidity affecting visibility [6, 7]. Standard VidLMs designed for terrestrial scenario environments struggle to perform effectively underwater due to these low-quality visual cues.
  • Lack of Scientific Domain Knowledge. Underwater content requires specialized ecological expertise for accurate interpretation. Unlike common scenarios featuring familiar objects and actions, underwater videos capture complex species interactions, specialized behavioral patterns, and environmental relationships that demand expert knowledge to decode [8, 9]. The interpretation of underwater footage requires understanding multiple layers of information, including taxonomic identification, morphological characteristics, behavioral states, and environmental contexts. Models must capture the intricate ecological relationships between marine organisms’ appearance, behaviors, and habitats, creating a significant knowledge gap for systems trained primarily on common objects and human activities.
  • Data Resource Scarcity. The development of effective video-language models for underwater environments is critically hindered by the absence of comprehensive training resources. Current video-language datasets predominantly focus on everyday human activities (HowTo100M [10]), sports (Sports-1M [11]), or general-purpose actions (Kinetics [12]), offering minimal support for specialized scientific domains. While underwater-specific datasets exist, they typically prioritize narrow tasks such as object tracking (WebUOT-1M [13]) or instance segmentation (UIIS [14]). These image-based datasets fail to equip models with the scientific domain knowledge necessary for comprehensive understanding of marine organism behaviors in underwater videos. This data gap creates a significant barrier to advancing video-language understanding beyond common scenarios to specialized scientific contexts.
To bridge these gaps, a comprehensive Underwater Video-Language Multimodal (UVLM) dataset integrates professional annotations is presented. Compared to existing video language understanding benchmarks, our UVLM demonstrates distinct differences in both content composition and construction methodology. First, in terms of content, UVLM is the first video-language benchmark specifically designed for underwater environments. To ensure the dataset accurately captures the distinctive characteristics of underwater settings, we carefully selected videos that encompass the unique challenges of this domain, including low-light conditions, water turbidity, and the highly variable movement patterns of marine organisms. Second, regarding construction methodology, we adopted a structured framework combining human-AI collaboration. The annotation targets encompass both biological and environmental elements, covering static and dynamic scenarios. After frame-by-frame manual annotation, we leveraged these human-annotated labels to guide GPT-4o in generating diverse sample content. Finally, all data underwent manual verification, with inconsistent entries either re-annotated or supplemented using search engines for factual accuracy.
The final dataset comprises 0.9M frames, 419 distinct biological categories, and diverse underwater scenarios, encompassing 20 types of video-language understanding tasks. Additionally, we established six specialized metrics specifically designed for quantitative performance comparison on the dataset. UVLM advances video-language research through its comprehensive coverage of core technical challenges:
  • Temporal Understanding: Supports development of models that can interpret continuous behavior sequences and environmental changes over time.
  • Fine-Grained Recognition: Enables research on distinguishing subtle visual differences with significant scientific meaning.
  • Compositional Reasoning: Facilitates the development of models that can decompose complex scenes into scientifically meaningful components.
  • Knowledge-Grounded Generation: Provides a foundation for generating technically accurate language descriptions based on visual evidence.
In summary, UVLM represents a significant step toward extending the capabilities of video-language models beyond everyday scenarios to specialized scientific domains. By providing richly annotated video-language pairs in underwater environments, UVLM enables the development of models that can interpret complex ecological dynamics and communicate this understanding through natural language, ultimately contributing to both multimedia research and marine science.
2 Related Work
Table 1: Representative video language benchmarks
Name Venue Themes & Scenarios
MSVD-QA [15], Next-QA [16] $M\prime$17, CVPR $\prime$21 Daily life
[…]
EgoSchema [18] NeurIPS $\prime$23 Life and human action, egocentric perspective
Video-Bench [19] arXiv $\prime$23 Television, music videos, NBA highlights
InternVid [20] ICLR $\prime$23 People and blogs, education, news and politics, how-to and style, entertainment, gaming, science, sports
MovieChat-1K [21] CVPR $\prime$24 Movie clips
MVBench [4] CVPR $\prime$24 Life, human action, movies, people, pets, science, instruction, finance, news, sports, humor, etc.
LONGVIDEOBENCH [22] NeurIPS $\prime$24 Life, movie, knowledge and news
E.T. Bench [23] NeurIPS $\prime$24 Indoor activities, tabletop, sports, egocentric views, cooking, news and vlogs, how-to, open-ended
LVBench [24] arXiv $\prime$24 Sports documentary, event record, lifestyle, TV show, cartoon
EgoThink [25] CVPR $\prime$24 Daily life understanding from first-person perspective
[…]
2.1 Video Language Understanding Benchmark
In recent years, many video language benchmarks have emerged, each targeting specific themes and application domains. Some focus primarily on everyday life scenarios  [15, 16, 17], while others emphasize human action or movie clips  [18, 21]. More comprehensive datasets cover a wider range of categories, including knowledge, sports, and instructional videos  [4, 22, 26, 27]. Table 1 provides a concise summary of representative benchmarks, highlighting their respective content coverage. Although these efforts have significantly advanced video-language research in terrestrial scenario contexts, the underwater environment remains largely uncharted, motivating our investigation of underwater video-language benchmark.
2.2 Underwater Observation Datasets
Underwater observation datasets and benchmarks have followed a clear progression, evolving from low-level tasks to high-level ones, from single-frame image analysis to video content analysis, and from unimodal to multimodal approaches. We summarize several key underwater image datasets and compare them with our UVLM in Table 2.
Low-level perception datasets focus on fundamental image enhancement and quality assessment, such as LSUI [28]. These datasets provide essential resources for developing and validating algorithms to address the typical degradation in underwater imagery. More recent efforts have extended this work into the temporal domain with video datasets like DRUVA [29], which contains 6,000 frames. Mid-level recognition tasks have also benefited from dedicated datasets. For instance, Wildfish [30] is tailored for marine life recognition, while COU [31] supports segmentation research. Additionally, a range of tracking datasets, such as VMAT [32] and WebUOT-1M [13] has further advanced object-centric underwater understanding. At the high-level, multimodal tasks are beginning to emerge. The MarineInst dataset [33] marks a critical step forward by supporting advanced tasks such as image segmentation and captioning, thereby opening new avenues for comprehensive underwater analysis.
Although these advancements have significantly propelled multimodal underwater analysis, multimodal underwater video analysis remains largely unexplored. Underwater videos inherently capture rich temporal dynamics, such as marine life trajectories and multi-view morphological changes, which provide additional contextual cues. By jointly leveraging visual appearance, temporal evolution, and domain-specific textual knowledge, unique advantages emerge. This holistic approach holds considerable promise for enhancing marine life interpretation, ecological monitoring, and overall environmental understanding.
Table 2: Comparison of underwater observation datasets.
Dataset Venue Im Vid Lang Seq Frame Task Categ.
[…]
UVLM - 2111 0.86M Video Understanding 419
2.3 Underwater Visual Analysis
Research efforts in this area [36] have focused primarily on three major directions: enhancement of underwater image quality [37, 38], interpretation of underwater objects [30, 31], estimation and modeling of the underwater environment [39, 40].
Underwater images typically have quality degradation caused by the scattering and absorption of light, color distortion, etc. To address these challenges, numerous underwater image enhancement algorithms [37, 41, 42, 38] have been proposed to restore visual information. Besides, for underwater perception, researchers have developed numerous specialized methods, including fish recognition [30], object detection [31], instance segmentation [43], and object tracking [13, 32], substantially advancing object-centric understanding in underwater environments.
Beyond the interpretation of objects, substantial progress has also been achieved in vision-based underwater environment estimation and modeling. Recent techniques such as dense 3D mapping [39], depth estimation [44, 45, 42] using diffusion models, and underwater scene reconstruction based on neural radiance fields [40] have greatly advanced our capabilities in comprehensive environmental understanding and reconstruction. Building upon the notable successes of single-modality underwater vision, Zheng et al. [33] further introduced multimodal approaches by incorporating textual information into underwater visual analysis.
In summary, underwater vision has seen steady progress, from low-level enhancement to high-level understanding. However, due to data limitation, the advanced VidLMs have yet to be effectively applied in underwater scenarios. We hope the proposed UVLM can help bridge this gap.
3 The proposed UVLM
Figure 2: An overview of data preparation and generation pipeline for UVLM.
3.1 Overview
Currently, research on VidLMs and the construction of corresponding benchmarks has been ongoing for several years, achieving significant progress. However, existing work primarily focuses on land scenarios, neglecting underwater environments. Extending VidLM technology to the underwater domain still faces challenges such as visual feature degradation caused by harsh aquatic conditions, the lack of domain knowledge due to domain gaps, and the scarcity of benchmarks due to difficulties in data acquisition.
In this paper, taking these challenges into consideration, we developed UVLM, the first underwater video-language benchmark. Firstly, we collect videos from typical underwater environments like oceans, lakes, and rivers, which exhibit different degradation problems such as water surface ripples, turbid water, and light scattering. Secondly, we select videos containing different content types, including static environmental features, static observations of organisms, dynamic scene variations, and dynamic biological behaviors, etc. To systematically inject domain knowledge into the dataset, we construct textual descriptions from 20 distinct perspectives covering observation content, behavioral analysis, and environmental dynamics. Thirdly, to maximize the scale of our dataset, we adopt a dual approach: systematically collecting data from online sources while strategically screening and re-annotating existing relevant datasets.
The final dataset comprises approximately two thousand carefully selected video sequences and 0.9 million frames, covering 419 different marine organisms and about 40 thousand video-language pairs. Previous underwater observation datasets [31, 43, 46, 47, 48, 49, 50, 51] typically focus on traditional computer vision tasks such as image segmentation and single object tracking, which are pure vision tasks. A few underwater observation datasets [33, 52] also incorporate visual and language elements, but the tasks are still limited to typical applications such as captioning. Unlike these datasets, the proposed UVLM is designed to incorporate specific underwater domain knowledge and requirements. To achieve this, we designed 16 to 20 relative questions referring to underwater research research topics such as marine organism recognition, behavioral analysis and prediction, habitat pattern characterization, subaquatic environmental monitoring, and ecosystem dynamics assessment, etc., covering both biological and environmental dimensions critical to underwater exploration and conservation. A comparison with previous underwater datasets is presented in Table 2.
3.2 Benchmark Construction
As shown in Figure 2, the benchmark construction consists for four major parts: video collection, manual annotation, model-assisted annotation and manual correction.
Video collection. Our objective extends beyond merely constructing an underwater video-language benchmark and training VidLMs on the data. Instead, we aim to build a dataset that captures the unique challenges of the underwater environment and enables VidLMs to operate in underwater contexts to support related research. To achieve this, video collection must satisfy two criteria: 1) the videos must capture distinctive underwater challenges; 2) the dataset must incorporate an adequate volume of representative cases to enable effective VidLM fine-tuning. To meet these requirements, we implement a dual-path video acquisition strategy.
The first path is collecting underwater videos from websites such as youtube and bilibili. We collect web-crawled data while enforcing two quality control principles during acquisition: 1) Both camera and observation objects or scenes must be under water. This selection rule is more in line with real world underwater vehicle and robot application scenarios; 2) The typical underwater scenarios and targets with different challenges and characteristics must be covered. The videos should cover the typical environments, like oceans, lakes, rivers and fish tanks, etc. Meanwhile, the chosen targets should include (but not limited to) common animals (fish, whales, prawns, tortoises, etc.), people, mermaids and generic objects. With the principles, we selected about 4 hundred video sequences, each of which contains 100 to 3000 frames, covering 53 categories of objects. The second path is re-annotating existing dataset. For scalability considerations, we selected WebUOT [13] as our primary data source.
Manual annotation. As emphasized in our introduction, the substantial domain gap between terrestrial scenario and underwater environments requires auxiliary annotation information to effectively leverage existing VidLMs for collaborative labeling. To facilitate granular analysis of underwater targets, such as studying the behavioral traits and ecological patterns of marine organisms, we annotate each frame with both bounding boxes and fine-grained taxonomic classifications aligned with marine biology standards. These annotations function as priors when fed into AI models for text generation.
For videos collected from websites, we employed 12 annotators annotate videos following three principles: 1) For occluded objects, only the visible part is marked with a rectangular box; 2) For the sharply protruding part, such as the tail and mouth of fish, whether to contain it in the bounding box is determined based on the proportion of the target and background in the additional introduced area. In specific, if the object’s protruding parts accounts for more than one-third of the additional introduced area, the sharply protruding part would be marked, otherwise it should not be marked; 3) Each frame is annotated with an axis aligned bounding box using the DarkLabel toolbox [^1] https://github.com/darkpgmr/DarkLabel, and each sequence has been assessed by three domain experts to minimize annotation errors. Unqualified labels would be sent to other annotators for re-labeling.
For videos from WebUOT, we followed several steps to refine video quantity. We first performed data cleaning to filter out videos with information interference, such as those containing large areas of subtitles or watermarks that affect video quality. Then, we conducted scene cleaning to remove videos that were not in natural underwater environments, such as those filmed in aquariums, fish tanks, or simulated gaming scenes. Next, we adopted SAM [53] and LaMa [54] to remove small-area watermarks, subtitles, and other information interference. We use mouse clicks to select the area, SAM for segmentation, and then LaMa for inpainting. Next, to avoid challenges caused by excessively long videos and focus specifically on evaluating the impacts from underwater environments, we split some overly long videos into segments of 300-600 frames each. Finally, based on statistical analysis, we removed some videos to balance the distribution of different observed object categories. In post-processing, a total of 544 videos were filtered out.
Model-assisted annotation. After preparing the videos, we carefully designed prompts to guide GPT-4o in generating relevant questions and answers based on the input video content. To ensure the generated questions have practical domain relevance, we first conducted research on key topics in marine biology, such as observed species, organism behavior, and habitat characteristics covered in the videos. Guided by these findings, we designed prompts to instruct GPT-4o to generate content along two dimensions: 1) Marine Organism Dimension. Static aspects: Species identification, biological attributes (e.g., morphology, coloration). Dynamic aspects: Behavioral analysis (e.g., feeding, interactions), movement patterns. 2) Underwater Environment Dimension. Static aspects: Environmental features (e.g., substrate type, coral structures), habitat traits. Dynamic aspects: Light condition variations, visibility fluctuations, etc.
The questions include two formats: multiple-choice and open-ended. For instance, multiple-choice: "What marine species is primarily observed in the images? A) Clownfish B) Chromis dimidiatus C) Parrotfish". Open-ended: "What is the overall setting of the video, and how does it influence Chromis dimidiatus’ activities?". During the Q&A process, we implemented clear prompt constraints, including: 1) Topic-specific constraints (e.g., focusing on taxonomic details, observed behaviors, or physicochemical properties of the water body); 2) Style requirements (e.g., diversifying sentence structures and enriching question types). Each video generates 16 to 20 video-text pairs. Please refer to the Appendix A for our generation prompts.
Manual correction. Despite the integration of domain experts and model-assisted generation in our annotation pipeline, the intrinsic complexity of underwater scenes and the limitations of automatic generation inevitably introduce semantic drift and consistency issues. To guarantee the ecological validity and scientific rigour of the final dataset, we therefore incorporated a dedicated manual-correction stage as the decisive quality-assurance step.
Stringent quality control was applied to all Q&A pairs generated by GPT-4o through a two-tier human review protocol. The first round, conducted by general reviewers, focused on detecting conflicts between the information supplied to GPT-4o and the content it produced. The second round involved senior experts who performed in-depth edits to ensure factual precision and domain conformity. During both rounds, hallucinations or statements irrelevant to the video context were either excised or fully rewritten. Whenever a model description diverged from the ground truth, we leveraged the accompanying assistant information to cross-validate and revise the erroneous content, thereby safeguarding the integrity of the dataset.
Figure 3: Statistics on UVLM. (a) Scene distribution; (b) Observation target distribution; (c) Object categories
3.3 UVLM Statistics
The scene distribution of UVLM and the distribution of various observation targets are presented in Figure 3. The final dataset comprises 2,109 videos (0.86M frames total), with lengths ranging from 100 to 3,000 frames. Spanning 419 categories across 4 major underwater scenes, each video contains 16–20 video-language pairs, yielding approximately 40k video-text pairs in total.
3.4 Evaluation Metric
Inspired by LLM judge [55] and MMDU [56], we build the evaluation pipeline by using GPT-4o. Specifically, based on GPT-4o, we evaluate the predictions from six aspects: 1) Semantic accuracy, assesses how closely the description matches the reference answer; 2) Visual perception accuracy, examines the correctness of the model’s interpretation of image content; 3) Detail completeness, measures the thoroughness of the provided information; 4) Environmental description accuracy, focuses on the correctness of underwater environmental descriptions; 5) Species behavior matching, evaluates the accuracy of prediction; 6) Choice accuracy, assesses the correctness of multiple-choice question. Please refer to the Appendix B for our evaluation prompts.
4 Experiments
Table 3: Performance comparison of different methods.
Method
Multiple
Choice
Accuracy
Semantic
Accuracy
Detail
Completeness
Visual
Perception
Accuracy
Environmental
Description
Accuracy
Species
Behavior
Matching
Overall
Closed-source VidLMs and other Open Source Large VidLMs
GPT4o 77.72 77.67 73.40 78.23 80.07 79.73 77.95
Claude3.7-Sonnet 76.61 73.35 73.58 74.10 79.71 76.35 76.09
Gemini2.5-Flash 78.22 72.43 73.34 70.53 78.32 74.92 75.00
Qwen2.5VL-72B 75.97 74.22 71.94 74.85 78.45 77.40 75.49
Base VidLMs
InternVL2.5-1B 48.54 44.35 47.30 45.30 50.15 44.85 46.73
VideoLLaMA3-2B 57.17 55.89 58.48 58.53 63.15 55.74 58.39
InternVL2.5-8B 57.64 57.70 59.75 59.45 63.48 61.05 60.15
VideoLLaMA3-7B 63.83 60.43 62.07 62.05 66.97 61.41 62.70
InternVL2.5-1B + UVLM 64.06 54.63 57.14 55.52 64.06 59.14 59.04
+15.52 +10.28 +9.76 +10.22 +13.91 +14.29 +12.31
VideoLLaMA3-2B + UVLM 71.96 63.66 65.35 65.74 72.48 67.70 67.61
+14.79 +7.77 +6.87 +7.21 +9.33 +11.96 +9.22
InternVL2.5-8B + UVLM 70.92 65.55 66.10 66.55 70.45 70.17 68.25
+13.28 +7.85 +6.35 +7.10 +6.97 +9.12 +8.10
VideoLLaMA3-7B + UVLM 76.92 70.52 70.17 71.39 76.41 73.39 73.03
+13.09 +10.09 +8.10 +9.34 +9.44 +11.98 +10.33
4.1 Dataset Partition
The proposed UVLM contains 2109 videos, covering 419 categories. As shown in Figure 3 (b), the overall category distribution follows a long-tail distribution and videos of many specific categories have relatively few observed targets. To ensure the training and test sets have consistent category distributions as much as possible, we sampled the test set proportionally. For categories with more than 5 videos, we randomly selected videos for the test set according to a predefined train-test split ratio. Then, we further randomly sampled videos from categories with fewer than 5 videos as test samples. Finally, 208 videos and their corresponding video-text pairs were selected as test set.
4.2 Results on UVLM
Table 3 and Figure 4 present the results on UVLM. We selected two recently released VidLMs, InternVL and Video-LLaMA, as baselines. For InternVL, we evaluated the 1B and 8B variants, while for Video-LLaMA, we tested both the 2B and 7B versions. According to the experimental results, we can draw several observations. First, underwater observation is indeed different from in-air scenarios. Even for current SOTA closed-source models like GPT-4o and Gemini, the scores only reach 77.95 and 75.00, respectively. Among open-source models, Qwen2.5VL-72B achieves a score of just 75.49. This indicates that there remains significant room for further improvement in underwater world understanding capabilities. Second, fine-tuning with UVLM significantly enhances VidLMs’ underwater observation capabilities. For instance, VideoLLaMA3-7B achieves a more than 10 points improvement in accuracy after fine-tuning, reaching an overall score of 73.03. At just 7B parameters, it trails Qwen2.5VL-72B (75.49) by only 2.46 percentage points, which demonstrates remarkable efficiency. Moreover, its performance becomes comparable to closed-source models like Gemini (75.00), despite the latter’s likely larger scale and proprietary training data. In fact, the benefits of UVLM fine-tuning extend beyond underwater observation, it also enhances performance on traditional benchmarks such as VideoMME and Perception Test. For detailed experimental results, please refer to the Appendix C.
5 Conclusion
In this paper, we introduced the first multimodal underwater observation benchmark specifically designed for enhancing video language model applications within underwater environments. Our benchmark integrates diverse and rich datasets, combining human expertise and advanced AI annotation methods to create an extensive resource suitable for multimodal instruction tuning. This dataset uniquely includes videos with varied resolutions, frame rates, and extensive marine biodiversity, covering 419 marine animal classes along with diverse static plants and terrains. Additionally, we meticulously addressed inherent underwater challenges, such as varying illumination, turbidity, and diverse viewing angles, providing realistic conditions for robust model training and evaluation. The benchmark incorporates comprehensive task diversity across 20 distinct observation scenarios, ranging from static content and dynamic behaviors to static and dynamic scene analyses. By proposing carefully designed and challenging evaluation metrics, our benchmark enables meaningful quantitative comparisons and facilitates the development of high-accuracy, reliable underwater observation systems. We anticipate this dataset will serve as a foundational resource for future research, encouraging innovation and promoting environmentally sustainable and economically viable solutions for underwater monitoring and exploration.
Figure 4: Visualization examples of fine-tuning with UVLM. Error descriptions are marked in red, and the accurate descriptions are marked in green.
6 Limitation & Broader Impacts
UVLM is the first large-scale underwater video–language benchmark, spanning 419 marine species, 20 task types. This rich coverage provides a solid foundation for robust, fair comparison of multimodal models in challenging aquatic conditions. While a small fraction of annotations may contain noise or class imbalance and extreme scenarios warrant further expansion, the present scale and quality already enable reliable training and analysis. Expert cross-checks, detailed documentation, and clear usage terms mitigate potential bias and misuse, while inviting the community to broaden geographic and modality coverage for improved generalization. UVLM is poised to accelerate marine-ecosystem monitoring, sustainable fisheries management, and public education, fostering cross-disciplinary collaboration and helping establish a secure, transparent, and responsible AI ecosystem for underwater research.